OpenStudy (anonymous):

I need help solving y^(prime) + 2xy = 2x^3. It's an ordinary differential equation.

OpenStudy (anonymous):

\[y \prime+2xy=2x^3\] \[I.F.=e ^{\int\limits 2xdx}=e ^{x^2}\] \[C.S. ~is~y~e ^{x^2}=\int\limits 2x^3 .e ^{x^2}dx=\int\limits~x^2~e ^{x^2}(2x)dx+c=I _{1}+c\] \[I _{1}=\int\limits x^2~e ^{x^2}(2x)dx\] put \[x^2=t,2x~dx=dt\] \[I _{1}=\int\limits t~e^t~dt\] integrate by parts and then replace the value of t.